OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (campbell_st):

for infinite solutions you need to look to see if 1 line is a multiple of the other... so looking at the last pair 5x - 3y =12 15x - 9y = 24 multiply the 1st equation by 3 15x - 9y = 36 so what has happened here is that the likes are parallel... same coefficients of x and y, but a different intercept...1st line has a y intercept y = 4 2nd line has a y intercept of y = 2 2/3 hope it helps